OpenStudy (anonymous):

Explain why the equation has no solution. square root(x-7) = 2

OpenStudy (anonymous):

it has a solution . replace x by 11 and it works

OpenStudy (anonymous):

x=11 is a solution....

OpenStudy (anonymous):

what chessmachine said

OpenStudy (anonymous):

e4

OpenStudy (anonymous):

c5

OpenStudy (anonymous):

\[\sqrt{x -3} = 2\]

OpenStudy (anonymous):

d4

OpenStudy (anonymous):

cxd4

OpenStudy (anonymous):

sorry, typed it wrong!

OpenStudy (anonymous):

c3

OpenStudy (anonymous):

dxc3 (EW, Smith Morra Gambit?)

OpenStudy (anonymous):

x=7 is still a solution....

OpenStudy (anonymous):

only way i know to throw people off their damned sicilian game. but since you know it i resign

OpenStudy (anonymous):

Lol gg

OpenStudy (anonymous):

i have beaten far superior players with this, just have to take advantage of the gain of tempo at some point before it is too late

OpenStudy (anonymous):

what about: \[\sqrt{x-3} = -2\]

OpenStudy (anonymous):

Well, GM Finegold of STL wrote a book about beating the Smith-Morra, and I've played him before ( I lived in STL for 3 years) and I was skimming his book at the STL Chess Club...so

OpenStudy (anonymous):

^^That one has no solution; there isn't a number you can plug in for x that will generate a negative.

OpenStudy (anonymous):

you played GM finegold then i resign before e4!

OpenStudy (anonymous):

It was a good (and equal) position about 20 moves in; I had the White pieces in a Caro-Kann main line, and then I made a horrible touch move blunder. T.T